OpenStudy (anonymous):

1) 5(x + 3) + 9 = 59 = Given 2) 5x + 15 + 9 = 59 = Distributive Property 3) 5x + 24 = 59 =? 4) 5x + 0 = 35 =? 5) 5x = 35 =? 6) x = 7 =? Match with one of these: A. Reflexive Property B. Symmetric Property C. Transitive Property D. Addition Property of Equality E. Subtraction Property of Equality F. Multiplication Property of Equality G. Division Property of Equality H. Substitution Property I. Additive Identity J. Multiplicative Identity K. Commutative Property L. Associative Property M. Distributive Property N. Given

OpenStudy (akashdeepdeb):

Can you elaborate the question?

OpenStudy (anonymous):

I need to find the property that each are so 1. 5(x + 3) + 9 = 59 I know is Given but I don't know the others

OpenStudy (anonymous):

I think it is 5x=59-24 x=35/5 x=7

OpenStudy (anonymous):

It could be substitution property